### v1

# Bayesian Sheaf Neural Networks

Abstract
Equipping graph neural networks with a convolution operation defined in terms of a cellular sheaf offers advantages for learning expressive representations of heterophilic graph data. The most flexible approach to constructing the sheaf is to learn it as part of the network as a function of the node features. However, this leaves the network potentially overly sensitive to the learned sheaf. As a counter-measure, we propose a variational approach to learning cellular sheaves within sheaf neural networks, yielding an architecture we refer to as a Bayesian sheaf neural network. As part of this work, we define a novel family of reparameterizable probability distributions on the rotation group $SO(n)$ using the Cayley transform. We evaluate the Bayesian sheaf neural network on several graph datasets, and show that our Bayesian sheaf models outperform deterministic sheaf models when training data is limited, and are less sensitive to the choice of hyperparameters.
Keywords: Cellular sheaf, Bayesian neural network, variational inference, Cayley transform
1 Introduction
Geometric and topological deep learning are rapidly developing research areas which, broadly speaking, aim to use geometric priors about the data of interest to develop neural network architectures using methods and tools from differential geometry, topology, and graph theory [12, 46]. Graph neural networks are a central example of geometric deep learning and have been applied to a wide range of tasks such as recommender systems [54], traffic modeling [42, 55], and predicting properties of chemicals [30]. Building on the basic idea of graph neural networks, architectures such as simplicial convolutional neural networks and hypergraph neural networks extend the graph convolution operation to a higher order convolution operation, which has advantages when the dataset contains salient relationships between groups of nodes rather than just pairs such as citation networks[20, 24], chemical data [8, 7], and neural data [44]. However, such convolution operators suffer from over-smoothing, and thus tend to perform best on homophilic graph data, where nodes connected together by edges or higher dimensional simplices have a high likelihood of belonging to the same class [50, 6].
Several approaches have been taken to address the limited capabilities of GCNs and higher-order variants for heterophilic graph data (i.e. not homophilic) [53, 48, 58, 57]. One approach replaces the graph (or simplicial) convolution operation with a convolution operation defined in terms of a cellular sheaf [32, 6]. A cellular sheaf, briefly, is a data structure defined on a simplicial complex which associates a vector space to each simplex, and a linear transformation known as a restriction map to each pair of incident simplices, that is, a pair of simplices where one simplex is a face of the other. The restriction maps provide additional flexibility in the message passing operation. They allow neighboring nodes to maintain distinct features throughout the layers of the sheaf neural network: an advantageous quality when working with heterophilic graph data.
While sheaf neural networks were developed to address the issue of over-smoothing in graph neural networks, they are still susceptible to many of the issues common to all neural networks, such as overfitting to limited training data and lack of robustness to noisy data, weight initializations, and hyperparameter selections. One approach to mitigating such issues is to apply tools and ideas from Bayesian statistics to deep learning [36, 47]. Bayesian approaches to machine learning have been used to quantify uncertainty of neural network predictions [38, 28], reduce overfitting [27, 29], add robustness to adversarial attacks [9], as well as simply improve the predictions of machine learning algorithms [22, 31, 43].
In this paper, we apply a Bayesian framework to learning cellular sheaves within a sheaf neural network. Specifically, we regard the sheaf Laplacian—an object which generalizes the graph Laplacian to cellular sheaves—as a latent random variable of the sheaf neural network.
Our main contributions are as follows.
  1. 1.
    We introduce Bayesian sheaf neural networks (BSNN) in which the sheaf Laplacian operator of the sheaf neural network is a latent random variable whose distribution is parameterized by a function of the input data. Motivated by the evidence lower bound for variational Bayesian inference, a pertinent loss function for our BSNN which contains a Kullback-Leibler (KL) divergence regularization term is derived to efficiently train the BSNN.
  2. 2.
    We define a novel family of reparameterizable probability distributions on the rotation group $SO(n)$ that we refer to as Cayley distributions. The Cayley distributions are reparameterizable and have tractable density functions, and thus can be used to perform variational inference on the sheaf Laplacian within the BSNN when the sheaf restriction maps belong to $SO(n)$. We derive closed-form expressions for the KL divergence of the Cayley distributions from the uniform distribution on $SO(n)$ when $n=2,3$.
  3. 3.
    We evaluate BSNN models on several benchmark graph datasets, specifically, on node-classification tasks with limited training data.
1.1 Related Work
We first discuss existing work on sheaf neural networks. Sheaf neural networks (SNN) were first introduced in [32], where they were shown to outperform graph convolutional networks on a semi-supervised node classification task using a hand-selected cellular sheaf. A more in-depth investigation of sheaf neural networks, sheaf diffusion processes, and the advantage of SNNs for heterophilic graph data was performed in [6], where they also constructed a method of learning a suitable cellular sheaf from the data as part of the sheaf neural network. Variants of sheaf neural networks include SNNs with attention-based edge weights [3], and SNNs defined for hypergraphs rather than cellular complexes [19]. Besides hand-crafting a sheaf for a particular problem or learning a sheaf during training, Barbero et al. [2] construct a sheaf at preprocessing time by assuming the graph data lies on a low dimensional manifold and defining a cellular sheaf in a manner inspired by connections from differential geometry.
Although no existing work has applied a Bayesian framework to sheaf neural networks, we mention some of the ways in which Bayesian neural networks have been combined with geometric machine learning. A variational auto-encoder using graph convolutional layers was considered in [40] for an edge detection task. Davidson et al. [18] discuss variational auto-encoders where the latent variable takes values in a hypersphere. Hajiramezanali et al. [31] combine graph variational autoencoders with graph recurrent networks to outperform existing state-of-the-art methods on dynamic link prediction tasks. Beyond variational autoencoders, Fan et al. model attention weights in graph attention networks as random latent variables, as well as in other architectures which use attention [22]. Given that the inclusion or exclusion of edges in graph datasets can be due to spurious noise or modelling assumptions, Zhang et al. view graph datasets as realizations of parametric families of random graphs, and develop a Bayesian graph convolutional network accordingly [56].
Defining a distribution over cellular sheaves with special orthogonal restriction maps requires a convenient family of distributions over $SO(n)$. The most well-known probability distributions on $SO(n)$ are the matrix Langevin (or von Mises-Fisher) distribution [45, 15] and the matrix Bingham distribution [14]. However such distributions are not easily reparameterizable, which makes incorporating them within neural networks trained via backpropagation difficult. Falorsi et al. [21] use the matrix exponential map to construct reparameterizable distributions on $SO(d)$, however, these distributions in general do not have tractable density functions. In this paper, we define a family of reparameterizable distributions on $SO(n)$ defined via the Cayley transform which we refer to as Cayley distributions. These are distinct from the Cayley distributions defined in [41], and we compare the two families of distributions in Section 4. We also mention [35], where a generalized version of the Cayley transform was used to simulate distributions on Stiefel and Grassmann manifolds.
1.2 Overview
In Section 2, we review preliminary topics such as cellular sheaves and sheaf Laplacians. We review how the sheaf Laplacian can be used to define a sheaf neural network, and the connection to a sheaf diffusion process. We prove a result regarding the linear separation power of a sheaf diffusion process for sheaves consisting of special orthogonal maps. In Section 3, we describe our variational approach to learning cellular sheaves within a sheaf neural network. In Section 4, we define a reparameterizable family of distributions on $SO(n)$, which are used to define the variational distribution for cellular sheaves with special orthogonal restriction maps. In Section 5, we evaluate our BSNN on several node classification tasks for heterophilic web-page datasets. We end with a final discussion, as well as potential directions for future work in Section 6.
2 Preliminaries
2.1 Cellular Sheaves
A $k$--simplex is the convex hull of $k+1$ affinely independent points in $\mathbb{R}^{n}$. Familiar examples include points, lines, triangles, and tetrahedrons, which are 00, $1$, $2$, and $3$--simplices respectively. A face $\tau$ of a simplex $\sigma$ is the simplex spanned by any subset of the points defining $\sigma$. In this case we denote $\tau\trianglelefteq\sigma$, and may say that $\tau$ is incident with $\sigma$. A simplicial complex $K$ is a collection of simplices in $\mathbb{R}^{n}$ such that for any simplex in $K$, all of its faces are also in $K$, and whenever two simplices $\sigma,\tau$ in $K$ have a non-empty intersection $\sigma\cap\tau$, their intersection is a face of both $\sigma$ and $\tau$.
A cellular sheaf $\mathcal{F}$ on a simplicial complex $K$ is a structure which assigns a vector space $\mathcal{F}(\sigma)$ to each simplex $\sigma\in K$, and assigns to each incident pair $\tau\trianglelefteq\sigma$ a linear transformation $\mathcal{F}_{\tau\trianglelefteq\sigma}:\mathcal{F}(\tau)\to\mathcal{F}(\sigma)$ such that $\mathcal{F}_{\sigma\trianglelefteq\sigma}:\mathcal{F}(\sigma)\to\mathcal{F}(\sigma)$ is the identity map for every $\sigma$, and with the compatibility condition that $\mathcal{F}_{\tau\trianglelefteq\rho}=\mathcal{F}_{\sigma\trianglelefteq\rho}\circ\mathcal{F}_{\tau\trianglelefteq\sigma}$ whenever $\tau\trianglelefteq\sigma\trianglelefteq\rho$. In keeping with the terminology from sheaf theory [11], the vector space $\mathcal{F}(\sigma)$ is referred to as the stalk of $\mathcal{F}$ over $\sigma$, and the transformations $\mathcal{F}_{\tau\trianglelefteq\sigma}$ are referred to as restriction maps. Cellular sheaves can be defined on cell complexes, a slight generalization of simplicial complexes, hence their name. For simplicity of presentation, we restrict our attention to simplicial complexes.
The motivation for considering cellular sheaves in the context of geometric deep learning is that given a cellular sheaf $\mathcal{F}$, there is a corresponding sheaf Laplacian $L_{\mathcal{F}}$. Just as how the message passing operation in many graph convolutional neural networks is defined in terms of the graph Laplacian $L=D-A$ [40], the sheaf Laplacian $L_{\mathcal{F}}$ can be used to define a message passing operation.
To begin defining the sheaf Laplacian for a cellular sheaf $\mathcal{F}$ over a finite simplicial complex $K$, assume a total order is placed on the vertices of $K$, so that each simplex $\sigma\in K$ may be written $\sigma=[v_{0},v_{1},\dots,v_{n}]$ with $v_{0}<v_{1}<\cdots<v_{n}$. Given a $k$--simplex $\sigma=[v_{0},v_{1},\dots,v_{k}]$, we denote its $(k-1)$--dimensional faces by $\partial_{i}(\sigma):=[v_{0},\dots,\hat{v}_{i},\dots,v_{k}]$ for $i\in\{0,\dots,k\}$, where $\hat{v}_{i}$ denotes the omission of vertex $v_{i}$.
An $n$--cochain of $(K,\mathcal{F})$ is an element $x\in C^{n}(K;\mathcal{F})$, where $C^{n}(K;\mathcal{F}):=\bigoplus_{\sigma\in K_{n}}\mathcal{F}(\sigma)$ denotes the direct sum of the vector spaces $\mathcal{F}(\sigma)$ over all $n$--dimensional simplices $\sigma\in K_{n}$. Given a cochain $x\in C^{n}(K;\mathcal{F})$ and an $n$--simplex $\sigma$, let $x_{\sigma}$ denote the value of $x$ at $\mathcal{F}(\sigma)$, i.e. $x_{\sigma}=\pi_{\sigma}(x)$ for the projection $\pi_{\sigma}:C^{n}(K;\mathcal{F})\to\mathcal{F}(\sigma)$. The $n$--th coboundary map $\delta_{n}:C^{n}(K;\mathcal{F})\to C^{n+1}(K;\mathcal{F})$ is the linear transformation given by defining the value of $\delta_{n}(x)$ at an $(n+1)$--simplex $\sigma$ to be
$\delta_{n}(x)_{\sigma}=\sum_{i=0}^{n+1}(-1)^{i}\mathcal{F}_{\partial_{i}\sigma\trianglelefteq\sigma}((x)_{\partial_{i}\sigma}).$ (1)
The $n$--th sheaf Laplacian of $(K,\mathcal{F})$ is the linear transformation $L_{\mathcal{F}}^{n}:C^{n}(K;\mathcal{F})\to C^{n}(K;\mathcal{F})$ defined by
$L^{n}_{\mathcal{F}}=\delta_{n-1}\delta_{n-1}^{*}+\delta_{n}^{*}\delta_{n},$ (2)
where $\delta_{n}^{*}:C^{n+1}(K;\mathcal{F})\to C^{n}(K;\mathcal{F})$ denotes the adjoint of $\delta_{n}$. In the case that $\mathcal{F}(\sigma)=\mathbb{R}$ for all $\sigma\in K$ and $\mathcal{F}_{\tau\trianglelefteq\sigma}$ is the identity map on $\mathbb{R}$ for all pairs $\tau\trianglelefteq\sigma$, the $n$--th sheaf Laplacian $L_{\mathcal{F}}^{n}$ is simply the combinatorial Laplacian of the simplicial complex $K$ [25]. For further information on cellular sheaves and sheaf Laplacians, we refer the interested reader to [33].
2.1.1 Cellular Sheaves over Graphs
Although the sheaf Laplacians $L_{\mathcal{F}}^{n}$ for multiple dimensions $n$ could be used as message passing operations within a sheaf neural network, due to the prevalence of graph-based data, one is primarily concerned with cellular sheaves over graphs, which are $1$--dimensional simplicial complexes, and the 00th sheaf Laplacian $L_{\mathcal{F}}^{0}$, denoted by $L_{\mathcal{F}}$.
Given a sheaf $\mathcal{F}$ over a graph $G$ with $n$ nodes, and assuming the stalk $\mathcal{F}(u)$ over each node $u\in V(G)$ has the same dimension $d$, the sheaf Laplacian, $L_{\mathcal{F}},$ may be represented as a block matrix of size $nd\times nd$. The off-diagonal blocks $L_{ij}\in\mathbb{R}^{d\times d}$ are $L_{ij}=-\mathcal{F}_{u_{j}\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{u_{i}\trianglelefteq e},$ whenever there exists an edge between the nodes $u_{i}$ and $u_{j}$, and $L_{ij}=0$ otherwise. The diagonal blocks are $L_{ii}=\sum_{u_{i}\trianglelefteq e}\mathcal{F}_{u_{i}\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{u_{i}\trianglelefteq e},$ where the sum is over all edges $e$ such that $u_{i}\trianglelefteq e$. If we let $D$ denote the $nd\times nd$ block diagonal matrix whose diagonal blocks are $L_{ii}$, and all other blocks are zero matrices, then the normalized 00th sheaf Laplacian $\Delta_{\mathcal{F}}$ defined in [32] is $\Delta_{\mathcal{F}}:=D^{-1/2}L_{\mathcal{F}}D^{-1/2}.$ Note $D$ is symmetric positive semi-definite, hence there exists a matrix $D^{1/2}$ such that $D^{1/2}D^{1/2}=D$. Moreover, we only consider sheaves such that the restriction maps $\mathcal{F}_{u\trianglelefteq e}$ are invertible, in which case $D$ is invertible, and so $D^{-1/2}$ is well-defined.
2.2 Sheaf Diffusion
Let $G$ be a graph with $n$ nodes and with $d$--dimensional node features, and let $\mathcal{F}$ be a cellular sheaf over $G$ with $d$--dimensional stalks. We represent the node features as an $nd$--dimensional vector $X\in\mathbb{R}^{nd}$ so that the normalized sheaf Laplacian $\Delta_{\mathcal{F}}$, as a $nd\times nd$ block matrix, updates $X$ by matrix-vector multiplication $\Delta_{\mathcal{F}}X$. Motivated by how the graph convolutional layers of [40] can be viewed as a discretization of a heat diffusion process $\tfrac{d}{dt}X(t)=-\Delta X(t)$ defined in terms of the normalized graph Laplacian $\Delta$, the authors of [6] consider a similar diffusion process which now uses the sheaf Laplacian:
$X(0)=X,\quad\tfrac{d}{dt}X(t)=-\Delta_{\mathcal{F}}X(t).$ (3)
Applying the Euler method with unit time-steps to approximate solutions to Eq. (3) yields
$X(t+1)=X(t)-\Delta_{\mathcal{F}}X(t).$ (4)
The authors of [6] use Eq. (4) to define sheaf diffusion layers as follows. Allowing for $f\in\mathbb{N}$ distinct feature channels, let $X\in\mathbb{R}^{nd\times f}$ be a feature matrix. Given a sheaf Laplacian operator $\Delta_{\mathcal{F}(t)}$ for layer $t$, a sheaf diffusion layer is defined by setting
$X_{t+1}=X_{t}-\sigma\big{(}\Delta_{\mathcal{F}(t)}(I_{n}\otimes W_{t,1})X_{t}W_{t,2}\big{)}.$ (5)
Here $W_{t,1}\in\mathbb{R}^{d\times d}$ and $W_{t,2}\in\mathbb{R}^{f\times f}$ are weight matrices for layer $t$, $\sigma$ is a non-linear activation function, $\otimes$ denotes a Kronecker product.
As a measure of the expressive power of sheaf layers based on Eq. (4), Bodnar et al. [6] study the ability of the sheaf diffusion process in Eq. (3) to linearly separate the features of nodes belonging to different classes.
Definition 2.1.
A class of sheaves $\mathcal{H}^{d}$ with $d$--dimensional stalks is said to have linear separation power over a class of labeled graphs $\mathcal{G}$ if for every $G\in\mathcal{G}$, there exists a sheaf $\mathcal{F}\in\mathcal{H}^{d}$ such that the sheaf diffusion process determined by $\mathcal{F}$ on $G$ linearly separates the features $X(t)$ of the classes of $G$ in the time limit $t\to\infty$, for almost all initial conditions $X(0)$.
As part of a sheaf learning mechanism within a sheaf neural network, a fixed class of sheaves $\mathcal{H}^{d}$ is parameterized, and the input data is mapped to a suitable sheaf $\mathcal{F}\in\mathcal{H}^{d}$ through this parameterization. For a node classification task with $C$ classes, the collection of sheaves $\mathcal{H}^{d}$ used would ideally have linear separation power over the collection of connected graphs with $C$ classes. While one could always take $\mathcal{H}^{d}$ to be as general as possible, it can be preferable to use a more restricted collection of sheaves, as this can reduce the number of model parameters involved in learning a cellular sheaf. For example, learning cellular sheaves whose restriction maps are diagonal or special orthogonal matrices involves fewer parameters than when the restriction maps are general linear maps. Recall that a special orthogonal matrix, i.e. element of $SO(n)$, is a matrix $Q$ such that $Q^{T}=Q^{-1}$ and $\det(Q)=1$.
It was shown that the class of sheaves $\mathcal{H}^{d}_{\operatorname{diag}}$ with invertible diagonal restriction maps has linear separation power over the collection $\mathcal{G}_{C}$ of connected graphs with $C$ classes whenever $C\leq d$ [6]. However, for $\mathcal{H}_{so}^{d}$, the class of sheaves with $d\times d$ special orthogonal restriction maps, [6, Proposition 13] only guarantees the existence of a $d$ such that $\mathcal{H}_{so}^{d}$ has linear separation power over $\mathcal{G}_{C}$ when $C\leq 8$.
Remark 2.2.
In [6], sheaves with orthogonal restriction maps, rather than special orthogonal, are considered. However, since the sheaf learning mechanism in their SNN is a continuous parametrization of the orthogonal group $O(n)$ by the connected space $\mathbb{R}^{m}$, it must have its image contained in a connected component of $O(n)$. Given that this implies that the restriction maps $\mathcal{F}_{u\trianglelefteq e}$ will either all have determinant $1$ or all have determinant $-1$, the maps $\mathcal{F}_{u\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{v\trianglelefteq e}$ used to define the sheaf Laplacian must always have determinant equal to $1$. Thus it is effectively equivalent to learning restriction maps belonging to $SO(n)$.
[…]
Proposition 2.3.
For any $d\geq 1$, if $\mathcal{G}$ denotes the class of connected graphs with $C\leq 7\lfloor\tfrac{d}{4}\rfloor$ classes, then $\mathcal{H}_{so}^{d}$ has linear separation power over $\mathcal{G}$.
For a proof see Appendix A.
Remark 2.4.
An extra consequence of Proposition 2.3 is that it also provides a new upper bound on the stalk dimension $d$ necessary for $\mathcal{H}^{d}_{gen}$, the collection of sheaves with $d\times d$ general linear restriction maps, to have linear separation power over the collection of connected graph $\mathcal{G}_{C}$ with $C$ classes. This is due to the simple fact that $\mathcal{H}^{d}_{so}\subseteq\mathcal{H}^{d}_{gen}$.
3 Bayesian Sheaf Neural Networks
We briefly establish some notation. Let $P$ be a distribution on a measurable space $\mathcal{X}$ with density $p$ with respect to some reference measure $\mu$. If $X$ is a random variable with distribution $P$, we write $X\sim P$. For a function $f:\mathcal{X}\to\mathbb{R}$, we denote by $\mathbb{E}_{P}[f(X)]$ the expectation of $f(X)$ with respect to the distribution $P$, that is, $\mathbb{E}_{P}[f(X)]=\int_{\mathcal{X}}p(x)f(x)\,\mu(dx)$. If $Q$ is another distribution over $\mathcal{X}$, and $Q$ has density $q$ with respect to $\mu$, the Kullback-Leibler divergence of $P$ from $Q$ is
$\textnormal{KL}(P\;\|\;Q)=\int_{\mathcal{X}}p(x)\log\Bigg{(}\frac{p(x)}{q(x)}\Bigg{)}\,\mu(dx).$
To simplify notation, we will denote a distribution by its density function. For instance, an expression of the form $X\sim p(X|Y)$ denotes that the random variable $X$ is distributed according to the conditional distribution of $X$ given $Y$.
3.1 Problem Statement
$X$$A$$\phi$$\theta$$\mathcal{F}$$y$$L$
Figure 1: Probability graphical model of the Bayesian sheaf neural network for a semi-supervised node classification problem. The adjacency matrix of the graph is $A$, and $\theta$, $\phi$ are weights of the network.
Consider a semi-supervised node classification problem, in which we have a graph dataset $G=(V,E)$ together with a node feature matrix $X$, a collection of observed node labels, $y$, and a collection of unobserved node labels $y^{*}$, all of which are random samples. Motivated by the expressive power of the sheaf diffusion process, we model the conditional probability $p_{\theta}(y|X)$ using a sheaf neural network whose layers are given by Eq. 5, with network parameters $\theta$ and a cellular sheaf $\mathcal{F}$.
Rather than take the learned cellular sheaf $\mathcal{F}$ (or sheaves) to be a deterministic function of $X$ as in [6], we instead regard $\mathcal{F}$ as a latent random variable of our model. See Figure 1 for a depiction of the probability graphical model. Note that the posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ will be intractable due to the fact that the labels $y$ are modeled as the output of a neural network. Hence, to do approximate posterior inference of $\mathcal{F}$ given the feature matrix, $X$, the observed node labels $y$, and model parameters $\theta$, we assume the intractable posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ belongs to a predetermined family of variational distributions $q_{\phi}(\mathcal{F}|X,y)$ parameterized by weights $\phi$.
We begin by noting that the log-likelihood $\log p_{\theta}(X,y)$ can be written
$\log p_{\theta}(X,y)=\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F}|X,y))+\mathcal{L}(\theta,\phi;X,y),$ (6)
where
$\mathcal{L}(\theta,\phi;X,y):=\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}\big{[}-\log q_{\phi}(\mathcal{F}|X,y)+\log p_{\theta}(\mathcal{F},X,y)\big{]}.$ (7)
The quantity $\mathcal{L}(\theta,\phi;X,y)$ is known as the evidence lower bound or ELBO, due to the fact that the KL divergence from one distribution to another is always non-negative, hence $\log p_{\theta}(X,y)\geq\mathcal{L}(\theta,\phi,X,y)$. Note from Eq. (6) that maximizing the ELBO with respect to the parameters $\theta$ and $\phi$ has the effect of maximizing the log-likelihood of the data, as well as minimizing the KL divergence of the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ from the true posterior distribution $p_{\theta}(\mathcal{F}|X,y)$. For this reason, the ELBO is commonly used as quantity to be optimized within variational Bayesian inference. See [5] for an overview of variational inference. In order to maximize $\mathcal{L}(\theta,\phi;X,y)$ with respect to the parameters $\theta,\phi$, we use the following equivalent expression:
$\mathcal{L}(\theta,\phi;X,y)=\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}\big{[}\log p_{\theta}(X,y|\mathcal{F})\big{]}-\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F})).$ (8)
In many cases, a closed-form expression can be derived for the KL-divergence term in the right hand side of Eq. (8), leaving only the gradient of the expectation term with respect to $\theta$ and $\phi$ to be estimated. See Appendix C for calculations of the KL divergence in certain cases. Similar to [39], we estimate $\nabla_{\phi}\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log p_{\theta}(X,y|\mathcal{F})]$ by making use of the so-called reparameterization trick. In our case, for a random cellular sheaf $\mathcal{F}$ distributed as $\mathcal{F}\sim q_{\phi}(\mathcal{F}|X,y)$, the reparameterization trick expresses $\mathcal{F}$ as a deterministic function $\mathcal{F}=g_{\phi}(\epsilon,X)$ of random noise $\epsilon\sim p(\epsilon)$. This allows us to rewrite the expectation with respect to $q_{\phi}(\mathcal{F}|X,y)$ as an expectation with respect to $p(\epsilon)$ which in turn allows the gradient to be moved within the expectation. That is,
$\nabla_{\phi}\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log p_{\theta}(X,y|\mathcal{F})]$ $=\nabla_{\phi}\mathbb{E}_{p(\epsilon)}[\log p_{\theta}(X,y|g_{\phi}(\epsilon,X))]$
$=\mathbb{E}_{p(\epsilon)}[\nabla_{\phi}\log p_{\theta}(X,y|g_{\phi}(\epsilon,X))]$
$\approx\frac{1}{K}\sum_{k=1}^{K}\nabla_{\phi}\log p_{\theta}(X,y|g_{\phi}(\epsilon^{(k)},X)),$
where $\epsilon^{(k)}$ are independent and identically distributed samples of $p(\epsilon)$ for $k\in\{1,2,\dots,K\}$.
Hence, when the KL divergence term of Eq. (8) can be computed analytically, if $\mathcal{F}^{(k)}=g_{\phi}(\epsilon^{(k)},X)$ for $\epsilon^{(k)}\sim p(\epsilon)$, our estimator is
~$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\,\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F})).$ (9)
Here, the weight $\lambda$ of the KL divergence term is included to combat issues with the KL divergence term vanishing during training, which we cyclically anneal during training in a manner similar to [26]. When the KL divergence term of Eq. (8) cannot be computed analytically, we may estimate it in a similar manner by noting that $\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F}))=\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log(q_{\phi}(\mathcal{F}|X,y)/p_{\theta}(\mathcal{F}))]$. In this case, the estimator, also denoted by ~$\widetilde{\mathcal{L}}_{\lambda}$, is given by
~$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\Big{(}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\log\frac{q_{\phi}(\mathcal{F}^{(k)}|X,y)}{p_{\theta}(\mathcal{F}^{(k)})}\Big{)},$ (10)
where again, $\mathcal{F}^{(k)}=g_{\phi}(\epsilon^{(k)},X)$ for $\epsilon^{(k)}\sim p(\epsilon)$.
3.2 Variational Sheaf Learner
We consider a variational sheaf learner in three different cases: where the learned sheaf $\mathcal{F}$ has restriction maps which are assumed to be (1) general linear, (2) special orthogonal, or (3) invertible diagonal. In all cases, we make the simplifying assumption that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ from the previous section factors as the product $\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$. While this assumption somewhat restricts the family of variational distributions, it makes computing the KL divergence term in the ELBO tractable. For further details, see Appendix C. We take each factor $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ to be a distribution $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ which depends on the type of restriction map used within the sheaf, where the distributional parameters $\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e}$ are a function of the node features $X$ and $\phi$ as follows. For each incident node-edge pair $u\trianglelefteq e$, if $u^{\prime}\trianglelefteq e$ denotes the other node incident with $e$, a multi-layer perceptron with learned weights $\phi$, denoted $\operatorname{MLP}_{\phi}$, maps the concatenated vector $[x_{u}||x_{u^{\prime}}]$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$:
$[\,\mu_{u\trianglelefteq e}\,||\,\sigma_{u\trianglelefteq e}\,]=\operatorname{MLP}_{\phi}([x_{u}||x_{u^{\prime}}]).$ (11)
Recall that a multi-layer perceptron is the composition of several feed-forward layers, each of which is a function of the form $x\mapsto\sigma(Wx+b)$, where $\sigma$ is a nonlinear activation function, in our case the exponential linear unit (ELU) [16], $W$ is a weight matrix, and $b$ is a vector of biases.
Case 1: General Linear. In the case where the sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are general linear maps, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d^{2}$--dimensional vectors. We take $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be normal with diagonal covariance so that for $z\sim\mathcal{N}_{d^{2}}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$ we set $\mathcal{F}_{u\trianglelefteq e}$ to be the vector $z$ rearranged into a $d\times d$ matrix. The reparameterization trick in this case writes $z=\mu_{u\trianglelefteq e}+\operatorname{diag}(\sigma_{u\trianglelefteq e})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d^{2}})$ so that $\mathcal{F}_{u\trianglelefteq e}$, being the matrix obtained by reshaping the vector $z$, is a deterministic function of $\mu_{u\trianglelefteq e}$, $\sigma_{u\trianglelefteq e}$, and $\epsilon$. We use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d^{2}}(\mathbf{0},I_{d^{2}})$.
Case 2: Special Orthogonal. The case where sheaf restriction maps are assumed to be special orthogonal presents challenges. Typical probability distribution on $SO(d)$, such as the matrix Langevin distribution, are not easily reparameterizable. This makes incorporating such distributions into neural networks, that are trained via back-propagation, difficult. On the other hand, while reparameterizable distributions on $SO(d)$ have been constructed in [21] using the exponential map, the density functions of such distributions can be difficult to compute exactly, which obstructs one from computing, or even efficiently estimating, the KL divergence term in Eq. (8). In Section 4 we define what we refer to as Cayley distributions, a family of reparameterizable distributions $\mathcal{C}_{d}(M,\kappa)$ on $SO(d)$ with closed form expressions for their probability density functions. We use $\mathcal{C}_{d}(M,\kappa)$ as the variational distributions for sheaves with special orthogonal restriction maps. Hence when $\mathcal{F}_{u\trianglelefteq e}\in SO(d)$, the distributional parameters are a mean rotation $\mu_{u\trianglelefteq e}\in SO(d)$ and scalar concentration parameter $\sigma_{u\trianglelefteq e}\in[0,1)$, and we have $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})=\mathcal{C}_{d}(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$. The reparameterization of the Cayley distributions is described in Remark 4.3. We use a prior distribution $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ where each $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ is a uniform distribution on $SO(d)$.
Case 3: Invertible Diagonal. In the case where the sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are assumed to be invertible diagonal matrices with respect to the standard basis, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d$--dimensional vectors and we take $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)=Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be a normal distribution with diagonal covariance, so that $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(z)$ for $z\sim\mathcal{N}_{d}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$. The reparameterization $\mathcal{F}=g_{\phi}(X,\epsilon)$ in this case is given by setting $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(\mu_{u\trianglelefteq e}+\sigma_{u\trianglelefteq e}\epsilon)$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d})$. Lastly, we use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d}(\mathbf{0},I_{d})$.
3.3 Bayesian Sheaf Neural Network Overview
Input Graph
MLP
Sheaf Conv.
Sheaf Conv.
$\cdots$
Sheaf Conv.
Sheaf Layers
Linear Layer
Output
[…]
Figure 2: Visual depiction of BSNN Architecture. The variational sheaf learning mechanism is depicted in the upper half of the diagram. Dashed arrows represent sampling from the sheaf distribution.
We give an overview of the Bayesian sheaf neural network architecture, which is also summarized in 2. The variational sheaf learning mechanism maps the input node features to the sheaf distributional parameters as described in Section 3.2, though we briefly describe here as well. A linear layer first resizes the input node features to dimension $k$. Specifically, we use $k=df$ where $d$ and $f$ are the stalk dimensions and number of feature channels respectively. Then, for each incident node-edge pair $u\trianglelefteq e$, the features of the pair of nodes corresponding to the edge $e$ are concatenated. Hence a node feature matrix of size $n\times k$ is transformed by this concatenation operation into a matrix of size $2e\times 2k$, where $n$ denotes the number of nodes, and $e$ denotes the number of edges in the graph. Lastly, an MLP maps these features indexed by the incident node-edge pairs $u\trianglelefteq e$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ as described by Eq. (11), which define the distribution over the cellular sheaves. From this distribution, a collection of independent and identically distributed sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$ are sampled: one for each layer of the network.
Separately, a multi-layer perceptron maps the input node features $X\in\mathbb{R}^{n\times m}$ to features of dimension $X^{\prime}\in\mathbb{R}^{n\times df}$ where $d$ is the stalk dimension and $f$ is the number of hidden channels. The features $X^{\prime}$ are reshaped to size $nd\times f$. Sheaf layers described by Eq. (5) update the transformed features $X^{\prime}$ using the sampled sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$. Finally, a linear layer maps the output of the final sheaf layer to a vector of class probabilities.
During training, a one-sample Monte Carlo estimate of Eq. (9) or Eq. (10) is computed, and the loss is back-propagated to update the network weights. During testing, ensembling is performed to obtain point estimates for the unobserved node labels $y^{*}$.
4 Reparameterizable Distributions on $SO(n)$
In order to perform variational inference for our BSNN when cellular sheaves with special orthogonal restriction maps are used, we require a family of distributions on $SO(n)$ that are both reparameterizable and have tractable density functions. We define a novel family of such distributions via the Cayley transform [13].
4.1 Cayley Distributions on $SO(n)$
We denote the set of all $n\times n$ skew-symmetric matrices by $\mathfrak{so}(n)$. The Cayley transform is a function $C:\mathfrak{so}(n)\to SO(n)$ defined by
$C(A)=(I-A)^{-1}(I+A)=2(I-A)^{-1}-I.$ (12)
If $A$ is an $n\times n$ skew-symmetric matrix, then $x^{\mathsf{T}}Ax=0$ for all $x\in\mathbb{R}^{n}$ due to the fact that $x^{\mathsf{T}}Ax=(x^{\mathsf{T}}Ax)^{\mathsf{T}}=-x^{\mathsf{T}}Ax$. Hence if $(I-A)x=0$, then $x^{\mathsf{T}}(I-A)x=x^{\mathsf{T}}x=0$ implying that $x=0$. Thus $I-A$ is invertible, showing that the Cayley transform is well-defined. If $P\in SO(n)$ does not have $-1$ as an eigenvalue, $C^{-1}(P)$ exists and is given by
$C^{-1}(P)=(P-I)(I+P)^{-1}=I-2(I+P)^{-1}.$ (13)
Hence $C$ is injective, and all but a measure zero subset of $SO(n)$ is contained in the image of $C$.
Remark 4.1.
The Cayley transform is sometimes alternatively defined as $C(A)=(I+A)^{-1}(I-A)=(I-A)(I+A)^{-1}$. Since the two definitions differ only by the precomposition of $C$ with the reflection $A\mapsto-A$, either choice of definition could be used without meaningful change.
For a vector $\phi\in\mathbb{R}^{n(n-1)/2}$, let $X_{\phi}$ denote the skew symmetric matrix containing the entries of $\phi$ as its lower triangular entries. Explicitly, if $\phi=(\phi_{1},\phi_{2},\dots,\phi_{n(n-1)/2})$ and if $x_{ij}$ denotes the $i,j$--th entry of $X_{\phi}$, then for $i>j$, set $x_{ij}=-x_{ji}=\phi_{ij-j(j+1)/2}$ and set $x_{ii}=0$. We may vectorize the Cayley transform and regard it as a map $\widetilde{C}:\mathbb{R}^{n(n-1)/2}\to\mathbb{R}^{n^{2}}$ by setting $\widetilde{C}(\phi)=\operatorname{vec}C(X_{\phi})$. The Jacobian (determinant) of $\widetilde{C}$ is $J\widetilde{C}(\phi)=\det(D\widetilde{C}(\phi)^{\mathsf{T}}D\widetilde{C}(\phi))^{1/2}$, where $D\widetilde{C}(\phi)$ denotes the total derivative of $\widetilde{C}$ at the point $\phi$. The Jacobian $J\widetilde{C}(\phi)$ can be expressed [41] in terms of $X_{\phi}$ by
$J\widetilde{C}(\phi)=\frac{2^{3n(n-1)/4}}{\det(I+X_{\phi})^{n-1}}.$ (14)
Definition 4.2.
Let $C:\mathfrak{so}(n)\to SO(n)$ denote the Cayley transform and let $X$ be a uniformly distributed random variable on $SO(n)$. Given parameters $M\in SO(n)$ and $10\leq\kappa<$ 1, we define the Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ on $SO(n)$ to be the distribution of $Y=C(\frac{1-\kappa}{1+\kappa}C^{-1}(X))M$.
Remark 4.3.
Note that by the definition of the Cayley distribution, a random variable $Y\sim\mathcal{C}_{n}(M,\kappa)$ can be expressed as a deterministic function of $M$, $\kappa$, and $X\sim U_{SO(n)}$. Hence these distributions are reparameterizable. For details on sampling from the uniform distribution on $SO(n)$, see for example [1].
In order to compute or efficiently estimate the KL divergence term in Eq. (9), an expression for the density of $\mathcal{C}_{n}(M,\kappa)$ is necessary. Using Eq. (14) for the Jacobian of the Cayley transform, we compute the density of $\mathcal{C}_{n}(M,\kappa)$ in the following theorem.
Theorem 4.4.
The Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ has density
$f_{n}(P;M,\kappa)=(1-\kappa^{2})^{\frac{n(n-1)}{2}}\det\Big{(}PM^{\mathsf{T}}-\kappa I\Big{)}^{1-n},$ (15)
with respect to the normalized Haar measure on $SO(n)$.
For a proof see Appendix B.
Remark 4.5.
The family of Cayley distributions defined in [41] have density proportional to $\det(PM^{\mathsf{T}}+I)^{\kappa}$ of Eq. (15). Note that the concentration parameter $\kappa$ appears in the exponent of the determinant expression, whereas in our proposed family of Cayley distributions, $\kappa$ instead scales the identity matrix term within the determinant.
4.2 Special Cases of the Cayley Distribution
We now examine the distributions $\mathcal{C}_{n}(M,\kappa)$ for small values of $n$.
Case $\mathbf{n=2}$. If $P,M\in SO(2)$ are rotations by $\theta,\mu$ radians respectively, then $\mathcal{C}_{2}(M,\kappa)$ has density
$f_{2}(P;M,\kappa)=\frac{1}{2\pi}\cdot\frac{1-\kappa^{2}}{1+\kappa^{2}-2\kappa\cos(\theta-\mu)}.$
This is the density of a wrapped Cauchy distribution on $SO(2)\cong S^{1}$, see[37]. Hence for the case $n=2$, the Cayley distributions coincide with known distributions.
Case $\mathbf{n=3}$. Note that $SO(3)$ is homeomorphic to $3$--dimensional real projective space $RP^{3}$, which is the quotient of $S^{3}$ under the identification of antipodal points. We show that $\mathcal{C}_{3}(M,\kappa)$ is related to an angular central Gaussian distribution [52] on $S^{3}$. Specifically, we show that $\mathcal{C}_{3}(M,\kappa)$ is the pushforward of an angular central Gaussian distribution on $S^{3}$ under the quotient map $\Phi:S^{3}\to SO(3)$. If $x\in\mathbb{R}^{n+1}$ is a unit vector, $\Lambda\in\mathbb{R}^{(n+1)\times(n+1)}$ is a symmetric positive definite matrix, and $\alpha_{n}$ denotes the volume of the $n$--dimensional sphere $S^{n}$, the angular central Gaussian distribution $ACG_{n}(\Lambda)$ on $S^{n}$ has density
$f(x;\Lambda)=\alpha_{n}^{-1}\det(\Lambda)^{-\frac{1}{2}}(x^{\mathsf{T}}\Lambda^{-1}x)^{-2}$
with respect to the uniform measure on $S^{n}$.
Proposition 4.6.
Let $\Phi:S^{3}\to\mathbb{R}\mathbb{P}^{3}\cong SO(3)$ denote the map which identifies antipodal points of $S^{3}$. For any $M\in SO(3)$ and $\kappa\in[0,1)$, there exists a symmetric positive definite matrix $\Lambda_{M,\kappa}\in\mathbb{R}^{4\times 4}$ such that if $X\sim ACG_{3}(\Lambda_{M,\kappa})$, then $\Phi(X)\sim\mathcal{C}_{3}(M,\kappa)$.
For a proof see Appendix B.
When $n=2$ or $n=3$, Eq. (14) for the Jacobian of the Cayley transform simplifies to $2^{3n(n-1)/4}(1+\|\phi\|^{2})^{1-n}$. One consequence of this fact is that $\mathcal{C}_{n}(I_{n},\kappa)$ is the pushforward of a Cauchy distribution (multivariate in the case $n=3$) under the Cayley transform. This fact makes it possible to compute the KL divergence of $\mathcal{C}_{n}(M,\kappa)$ from the uniform distribution on $SO(n)$ in these cases.
Proposition 4.7.
For $n=2,3$ the KL divergence of $\mathcal{C}_{n}(M,\kappa)$ from the uniform distribution is given by
$\textnormal{KL}(\mathcal{C}_{2}(M,\kappa)\;\|\;U_{SO(2)})$ $=-\log(1-\kappa^{2})$
$\textnormal{KL}(\mathcal{C}_{3}(M,\kappa)\;\|\;U_{SO(3)})$ $=-\log(1-\kappa^{2})-2\log(1-\kappa)-2\kappa.$
For a proof see Appendix C.
Thus when learning a distribution of cellular sheaves with special orthogonal restriction maps in our BSNN, for the cases where the stalk dimension $d$ is $2$ or $3$, we are able to use Eq. (9) to estimate the ELBO by plugging in the expressions for the KL divergence from Proposition 4.7. When $d\geq 4$, we estimate the ELBO by simply using Eq. (10).
5 Node Classification Experiments
We compare our Bayesian sheaf neural networks (BSNN) against the sheaf neural networks (NSD) from [6] on the WebKB datasets, a set of real-world web-page datasets. These datasets have low edge-homophily coefficients, that is, there are relatively few edges which connect nodes belonging to the same class [58]. This is a type of graph data where sheaf neural networks are expected to outperform traditional graph neural networks. See Appendix D for further details on the datasets. To study the performance of our Bayesian sheaf models when training data is limited, we use 10 fixed splits of the datasets, each of which contains $32\%/20\%/48\%$ of the nodes for training, validation, and testing respectively. For a given hyperparameter configuration, each model is trained and evaluated on 10 fixed splits of the data and we compute the average validation accuracy. Identical hyperparameter sweeps are performed for each model. For the hyperparameter configuration that yields the best validation accuracy for a model, we train and evaluate the model on the 10 splits for 30 different random seeds. We report the mean test accuracies along with the standard deviation of the test accuracies across the 30 random seeds in Table 1. For each type of restriction map (diagonal, special orthogonal, and general linear), we compare the results of the BSNN model against the NSD model. The higher test accuracy is bolded, and further, a blue entry denotes that a Wilcoxon signed rank test indicates a statistically significant difference for a $p$-value threshold of $p=0.05$. We also include results for a graph convolutional network (GCN) as a baseline comparison.
Additionally, to examine the sensitivity of the Bayesian and deterministic sheaf models to the hyperparameter selections, we report the mean test accuracy across all hyperparameter configurations in Table 2.
5.1 Results
Though the results are mixed in the case of diagonal restriction maps, the $SO(d)$-BSNN outperforms the corresponding $O(d)$-NSD model on all three datasets, and the Gen-BSNN significantly outperforms the corresponding Gen-NSD on all three dataset. This suggests that the comparative advantage of our Bayesian approach to learning cellular sheaves is more pronounced for sheaves with a greater number of parameters. We also note that on each of the three WebKB datasets, the best test accuracy out of all the models is attained by one of our BSNNs. Moreover, the standard deviation of the reported test accuracies for each BSNN is lower than the corresponding NSD model with the same type of restriction map in all but one case, showing that our BSNN models are less sensitive to the effects of random weight initializations.
From Table 2, the advantage of each BSNN model over the corresponding NSD model when considering the entire range of hyperparameters is definitive. For every type of restriction map and dataset, a Wilcoxon signed-rank test for a $p$-value threshold of $p=0.05$ indicates a statistically significant difference between the results of the BSNN model and the corresponding NSD model. Thus we observe that the BSNN models maintain good performance over a much wider range of hyperparameters when compared with the NSD model with the same type of restriction maps.
Table 1: Average test accuracies and standard deviations across 30 random seeds for the hyperparameter configuration yielding the highest validation accuracy for each model and dataset.
Texas Wisconsin Cornell
Diag-BSNN $75.35{\scriptstyle\pm 1.68}$ $\mathbf{83.04}{\scriptstyle\pm 0.87}$ $72.23{\scriptstyle\pm 1.76}$
Diag-NSD $\mathbf{75.49}{\scriptstyle\pm 2.40}$ $82.77{\scriptstyle\pm 1.15}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.70}}{\scriptstyle\pm 1.45}$
$\mathbf{SO(d)}$-BSNN $\mathbf{76.65}{\scriptstyle\pm 1.48}$ $\mathbf{82.02}{\scriptstyle\pm{1.06}}$ $\mathbf{74.31}{\scriptstyle\pm{1.10}}$
$O(d)$-NSD $76.09{\scriptstyle\pm 1.79}$ $81.79{\scriptstyle\pm 1.63}$ $74.07{\scriptstyle\pm 1.58}$
Gen-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}76.32}}{\scriptstyle\pm 1.65}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}82.03}}{\scriptstyle\pm{0.94}}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}72.29}}{\scriptstyle\pm{1.21}}$
Gen-NSD $72.97{\scriptstyle\pm 1.81}$ $77.25{\scriptstyle\pm 1.67}$ $70.16{\scriptstyle\pm 2.08}$
GCN $56.50{\scriptstyle\pm 0.81}$ $55.63{\scriptstyle\pm 1.13}$ $51.97{\scriptstyle\pm 0.49}$
Table 2: Average test accuracy across all hyperparameter configurations and corresponding standard deviations.
Texas Wisconsin Cornell
Diag-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.41}}{\scriptstyle\pm 2.84}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}77.78}}{\scriptstyle\pm 3.32}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}64.77}}{\scriptstyle\pm 5.51}$
Diag-NSD $71.48{\scriptstyle\pm 3.54}$ $74.85{\scriptstyle\pm 4.57}$ $62.18{\scriptstyle\pm 7.10}$
$\mathbf{SO(d)}$-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}72.88}}{\scriptstyle\pm 3.63}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}76.84}}{\scriptstyle\pm 4.68}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.21}}{\scriptstyle\pm 6.84}$
$O(d)$-NSD $71.98{\scriptstyle\pm 3.28}$ $74.55{\scriptstyle\pm 5.03}$ $61.73{\scriptstyle\pm 6.70}$
Gen-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.12}}{\scriptstyle\pm 2.41}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}77.92}}{\scriptstyle\pm 3.23}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.77}}{\scriptstyle\pm 4.87}$
Gen-NSD $71.34{\scriptstyle\pm 2.70}$ $74.64{\scriptstyle\pm 3.92}$ $62.20{\scriptstyle\pm 5.69}$
6 Conclusion and Future Directions
In this work, we have introduced Bayesian sheaf neural networks: novel sheaf neural networks in which the sheaf Laplacian is a random latent variable, and where the loss function used in training contains a Kullback-Leibler divergence regularization term. Not only does this allow one to quantify uncertainty of the model arising from the learned sheaf, but we show that for node classifications tasks with limited training data, our BSNN models outperform deterministic sheaf models. Additionally, our experiments suggest that the BSNN models are more robust to different choices of hyperparameters and weight initializations than the corresponding deterministic sheaf models.
We hope that this paper encourages further work on stochastic sheaf neural networks. We briefly mention some possible future directions. Different priors for the distribution of sheaf Laplacians could be suitable for the BSNN depending on the graph dataset, in particular, depending on the level of edge homophily. Although we view the underlying graph as fixed in this paper, one could view it probabilistically and combine our work with the approach taken in [56]. Lastly, similar to the continuous-time version of the sheaf diffusion network that was studied in [6], a continuous-time version of the BSNN could be considered, where the sheaf Laplacian is a stochastic process.
Acknowledgements
This work has been partially supported by the Army Research Laboratory Cooperative Agreement No. W911NF2120186, and STRONG ARL CA No W911NF-22-2-0139.
Appendix A Proof of Proposition 2.3
In what follows, given a cellular sheaf $\mathcal{F}$ over a graph $G=(V,E)$ and initial conditions $X(0)$, let $X(t)$ denote the solution to Eq. (3). Note that the limit of $X(t)$ as $t\to\infty$ exists, and is equal to the orthogonal projection of $X(0)$ onto $\ker(\Delta_{\mathcal{F}})$ [34]. If $\mathcal{F}$ has $d$--dimensional stalks and $G$ has $n$ nodes, although $X(t)\in\mathbb{R}^{nd}$, it will be useful to equally regard $X(t)$ as an $n\times d$ matrix. With this in mind, let $X(t)_{u}$ denote the row of $X(t)$ corresponding to node $u$. We write $u\in V_{l}$ if node $u$ has label $l$. For a cellular sheaf $\mathcal{F}$ with $d$--dimensional stalks over a graph $G=(V,E)$ with labeling $\ell:V\to\mathcal{L}$ and initial conditions $X(0)$, we say that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ if for each label $l\in\mathcal{L}$, the sets $\{\lim_{t\to\infty}X(t)_{u}:u\in V_{l}\}\subseteq\mathbb{R}^{d}$ and $\{\lim_{t\to\infty}X(t)_{u}:u\notin V_{l}\}\subseteq\mathbb{R}^{d}$ are linearly separable. With this terminology, Definition 2.1 states that a collection of cellular sheaves $\mathcal{H}$ has linear separation power over a collection of labelled graphs $\mathcal{G}$, if for each graph $G=(V,E)\in\mathcal{G}$ with labelling $\ell:V\to\mathcal{L}$, there exists a cellular sheaf $\mathcal{F}\in\mathcal{H}$ over $G$ such that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ for almost all initial conditions $X(0)$, i.e. all but a measure zero subset of the initial conditions.
Proof of Proposition 2.3.
Let $d\geq 1$ and let $d=4k+m$ for $40\leq m<$ 4. That is, $k=\lfloor\tfrac{d}{4}\rfloor$. Since proving the proposition for $C=7k$ also proves the cases $C<7k$, we may assume that $C=7k$. Let =$\mathcal{L}=\{1,2,\dots,C\}$ denote the set of class labels for the nodes of a graph $G=(V,E)\in\mathcal{G}$ and let $\ell:V\to\mathcal{L}$ denote the labelling of the nodes of $G$. For each $1\leq i\leq k$, let $J_{i}=\{7(i-1)+1,7(i-1)+2,\dots,7(i-1)+7\}$, let $\mathcal{L}_{i}=\{j:j\in J_{i}\}\cup\{0\}$, and let $\ell_{i}:V\to\mathcal{L}_{i}$ denote a relabeling of $G$ where for each $u\in V$, $f_{i}(u)=f(u)$ whenever $f(u)\in J_{i}$, and $f_{i}(u)=0$ otherwise. That is, the relabeling $\ell_{i}$ retains the seven labels in $J_{i}$, and replaces all other labels with a single distinct eighth label.
By [6, Proposition 13], for each $1\leq i\leq k$ we may find a cellular sheaf $\mathcal{F}^{i}\in\mathcal{H}_{so}^{4}$ with $4$--dimensional stalks which linearly separates the eight classes of $(G,X(0),\ell_{i})$ for almost all initial conditions $X(0)\in\mathbb{R}^{4n}$. Let $\mathcal{F}$ denote the cellular sheaf with restriction maps $\mathcal{F}_{u\trianglelefteq e}=\mathcal{F}^{1}_{u\trianglelefteq e}\oplus\mathcal{F}^{2}_{u\trianglelefteq e}\oplus\cdots\oplus\mathcal{F}^{k}_{u\trianglelefteq e}\oplus I_{m}$. We claim that $\mathcal{F}$ linearly separates the classes of the original labelling $(G,X(0),\ell)$ for almost all initial conditions $X(0)\in\mathbb{R}^{n\times d}$.
Given initial conditions $X(0)\in\mathbb{R}^{n\times d}$, for each $1\leq i\leq k$ let $X_{i}(0)\in\mathbb{R}^{n\times 4}$ be the matrix consisting of columns $34i-$ 3 through $4i$ of $X(0)$. Note that for almost all initial conditions $X(0)$, it will be true that for all $1\leq i\leq k$, $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$. Next, for any class $l\in\mathcal{L}$, find $i$ such that $l\in J_{i}$. Since $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$, we may find a hyperplane $A\subseteq\mathbb{R}^{4}$ which separates the class $l$ from all other classes in $\mathcal{L}_{i}$, and hence $\mathcal{L}$ as well. Then $\mathbb{R}^{4}\oplus\dots\oplus A\oplus\dots\oplus\mathbb{R}^{4}\oplus\mathbb{R}^{m}$ is a hyperplane in $\mathbb{R}^{d}$ separating class $l$ from all other classes. ∎
Appendix B Section 4 Proofs
For a subset $S$ of a metric space $M$ and for $\delta>0$, let
$\mathcal{H}^{d}_{\delta}(S)=\inf\Big{\{}\sum_{i=1}^{\infty}\alpha_{d}(\tfrac{1}{2}\operatorname{diam}(S))^{d}:\bigcup_{i=1}^{\infty}U_{i}\supseteq S,\operatorname{diam}(U_{i})<\delta\Big{\}}$
where $\alpha_{d}$ denotes the volume of a unit $d$--ball. The $d$--dimensional Hausdorff measure on $M$ is then given by $\mathcal{H}^{d}(S)=\lim_{\delta\to 0}H^{d}_{\delta}(S)$.
Let $\lambda^{n}$ denote the $n$--dimensional Lebesgue measure. The area formula [23] computes the Hausdorff measure of the image of a $1$-to-$1$ and continuously differentiable function $f$ in terms of the Jacobian of $f$. We state a version presented in [51].
Proposition B.1 (Area formula).
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be $1$-to-$1$ and continuously differentiable. For any Borel subset $A\subseteq\mathbb{R}^{n}$,
$\int_{A}Jf(x)\lambda^{n}(dx)=\int_{f(A)}\mathcal{H}^{n}(dy).$
Moreover,
$\int_{A}g(f(x))Jf(x)\lambda^{n}(dx)=\int_{f(A)}g(y)\mathcal{H}^{n}(dy),$ (16)
for all Borel measurable $g:\mathbb{R}^{m}\to\mathbb{R}$ for which one side of (16) exists.
Equip $SO(n)$ with the metric induced by the Frobenius norm $\|A\|=\operatorname{tr}(A^{\mathsf{T}}A)$. Since left (or right) multiplication on $SO(n)$ is an isometry, it follows that the Hausdorff measure is a left (and right) translation invariant measure on $SO(n)$. Moreover, if $d=n(n-1)/2$ denotes the dimension of $SO(n)$, then $\mathcal{H}^{d}$ is a locally finite measure on $SO(n)$, and thus is a Haar measure by Haar’s theorem.
Hence, Proposition B.1 yields a change of variables formula for they Cayley transform. For example, if $X$ is a uniformly distributed random variable on $SO(n)$, then $C^{-1}(X)$ has density $p(x)=J\widetilde{C}(x)$ with respect to the Lebesgue measure on $\mathbb{R}^{d}$.
Proof of Theorem 4.4.
[…]
$f_{n}(P;I,\kappa)=\gamma^{\frac{n(n-1)}{2}}J\widetilde{C}\Big{(}\gamma\widetilde{C}^{-1}(\operatorname{vec}P)\Big{)}J\widetilde{C}^{-1}(\operatorname{vec}P).$
Using Eq. (14), we obtain
$f_{n}(P;I,\kappa)=\gamma^{\frac{n(n-1)}{2}}\Big{(}\frac{\det(I+C^{-1}(P))}{\det(I+\gamma C^{-1}(P))}\Big{)}^{n-1}.$ (17)
Using the equality $\det(I+C^{-1}(P))=\frac{2^{n}}{\det(I+P)}$,
$f_{n}(P;I,\kappa)=2^{n(n-1)}\gamma^{\frac{n(n-1)}{2}}\det\Big{(}(I+P)(I+\gamma C^{-1}(P))\Big{)}^{1-n}.$
Next, from Eq. (13) we have that $I+\gamma C^{-1}(P)=(1+\gamma)I-2\gamma(I+P)^{-1}$. By substituting and simplifying,
$f_{n}(P;I,\kappa)=2^{n(n-1)}\gamma^{\frac{n(n-1)}{2}}\det\Big{(}(1+\gamma)(I+P)-2\gamma I\Big{)}^{1-n}.$
Factoring $1+\gamma$ out of the expression within the determinant, we have
$f_{n}(P;I,\kappa)=2^{n(n-1)}\Big{(}\frac{\gamma}{(1+\gamma)^{2}}\Big{)}^{\frac{n(n-1)}{2}}\det\Big{(}P-\frac{\gamma-1}{\gamma+1}I\Big{)}^{1-n},$
or in terms of $\kappa$,
$f_{n}(P;I,\kappa)=(1-\kappa^{2})^{\frac{n(n-1)}{2}}\det\Big{(}P-\kappa I\Big{)}^{1-n}.$
Proof of Proposition 4.6.
Representing points of $S^{3}$ as unit vectors $x=(a,b,c,d)\in\mathbb{R}^{4}$, the $2$--fold cover $\Phi:S^{3}\to SO(3)$ has an explicit description as
$\Phi(x)=\begin{pmatrix}a^{2}+b^{2}-c^{2}-d^{2}&2(bc-ad)&2(bd+ac)\\ 2(bc+ad)&a^{2}-b^{2}+c^{2}-d^{2}&2(cd-ab)\\ 2(bd-ac)&2(cd+ab)&a^{2}-b^{2}-c^{2}+d^{2}\end{pmatrix}.$
A straightforward calculation shows that
$\det(\Phi(x)-\kappa I)=(1-\kappa)^{3}a^{2}+(1+\kappa)^{2}(1-\kappa)(b^{2}+c^{2}+d^{2}).$
If $f(P;M,\kappa)$ denotes the density of $\mathcal{C}_{3}(M,\kappa)$ with respect to the unit Haar measure on $SO(3)$, $g(x;\Lambda)$ denotes the density of $ACG_{3}(\Lambda)$ with respect to the unit Haar measure on $S^{3}$, and $\Lambda_{\kappa}$ is the diagonal matrix $\Lambda_{\kappa}=\operatorname{diag}((\frac{1+\kappa}{1-\kappa})^{2},1,1,1)$, then we have
$f(\Phi(x);I,\kappa)$ $\propto\det(\Phi(x)-\kappa I)^{-2}$
$=\Big{(}(1-\kappa)^{3}a^{2}+(1+\kappa)^{2}(1-\kappa)(b^{2}+c^{2}+d^{2})\Big{)}^{-2}$
$\propto\Big{(}\big{(}\frac{1-\kappa}{1+\kappa}\big{)}^{2}a^{2}+b^{2}+c^{2}+d^{2}\Big{)}^{-2}$
$\propto g(x;\Lambda_{\kappa}),$
where the symbol $\propto$ denotes "proportional to".
Moreover, for $M\in SO(3)$, let $y_{M}\in\Phi^{-1}(M^{\mathsf{T}})$. Regarding $y_{M}$ as a unit quaternion, let $Q\in SO(4)$ be the rotation $z\mapsto zy_{M}$ for $z\in\mathbb{H}$. Note that $\Phi(x)M^{\mathsf{T}}=\Phi(xy_{M})=\Phi(Qx)$ and $g(x;Q^{\mathsf{T}}\Lambda_{\kappa}Q)=g(Qx;\Lambda_{\kappa})$. Hence
$f(\Phi(x);M,\kappa)=f(\Phi(Qx);I,\kappa)\propto g(x;Q^{\mathsf{T}}\Lambda_{\kappa}Q).$
Lastly, since $\Phi$ is a homomorphism, the pushforward of the unit Haar measure on $S^{3}$ yields the unit Haar measure on $SO(3)$. This combined with the fact that $f(\Phi(x);M,\kappa)\propto g(x;Q^{\mathsf{T}}\Lambda_{\kappa}Q)$ implies that $\Phi(X)\sim\mathcal{C}_{3}(M,\kappa)$ whenever $X\sim ACG_{3}(\Lambda_{M,\kappa})$ for $\Lambda_{M,\kappa}=Q^{\mathsf{T}}\Lambda_{\kappa}Q$. ∎
Appendix C KL Divergence Calculations
Recall that we assume that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ factors as $q_{\phi}(\mathcal{F}|X,y)=\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$, and that we use a prior $p_{\theta}(\mathcal{F})$ which also factors in the same manner $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$. As a consequence of the chain rule for KL divergence, see for example [17, Theorem 2.5.3], we have
$\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F}))=\sum_{u\trianglelefteq e}\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e})).$
Hence it suffices to be able to compute $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ for all $u\trianglelefteq e$.
In the cases where we learn diagonal or general linear sheaf restriction maps, $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ is a KL divergence of a normal distribution with diagonal covariance from a standard normal distribution. The KL divergence of a normal distribution $\mathcal{N}(\mu,\operatorname{diag}(\sigma^{2}))$ with mean $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}^{d}$ from the standard normal distribution is
$\textnormal{KL}(\mathcal{N}(\mu,\sigma^{2})\;\|\;\mathcal{N}(0,1))=-\frac{1}{2}\sum_{i=1}^{d}\big{(}1+\log(\sigma^{2}_{i})-\mu_{i}^{2}-\sigma_{i}^{2}\big{)}.$ (18)
In the case where the restriction maps are special orthogonal, and the stalk dimension $d$ is $2$ or $3$, we compute $\textnormal{KL}(q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)\;\|\;p_{\theta}(\mathcal{F}_{u\trianglelefteq e}))$ using Proposition 4.7, which we now prove.
Proof of Proposition 4.7.
First, we note that the integral of a function $f(P)$ on $SO(n)$ with respect to a Haar measure is invariant under the change of variables $P\mapsto PM$ for any $M\in SO(n)$. Hence $\textnormal{KL}(\mathcal{C}_{n}(M,\kappa)\;\|\;U_{SO(n)})$ does not depend on $M$, that is,
$\textnormal{KL}(\mathcal{C}_{n}(M,\kappa)\;\|\;U_{SO(n)})=\textnormal{KL}(\mathcal{C}_{n}(I,\kappa)\;\|\;U_{SO(n)}).$
To simplify notation, set $d=n(n-1)/2$, let $\gamma=\frac{1+\kappa}{1-\kappa}$, and let $V_{n}$ denote the volume of $SO(n)$. Recall Eq. (17), and note that the density of $\mathcal{C}_{n}(I,\kappa)$ with respect to $\mathcal{H}^{d}$ can be expressed as
$V_{n}^{-1}\gamma^{d}\Big{(}\frac{\det(I+C^{-1}(P)}{\det(I+\gamma C^{-1}(P))}\Big{)}^{n-1}$
where $V_{n}$ denotes the volume of $SO(n)$. Then the change of variables formula yields
$\textnormal{KL}(\mathcal{C}_{n}(M,\kappa)\;\|\;U_{SO(n)})=\int_{\mathbb{R}^{d}}\frac{2^{\frac{3n(n-1)}{4}}V_{n}^{-1}\gamma^{d}}{\det(I+\gamma X_{\phi})^{n-1}}\log\Big{(}\frac{\gamma^{d}\det(I+X_{\phi})^{n-1}}{\det(I+\gamma X_{\phi})^{n-1}}\Big{)}\,\lambda^{d}(d\phi),$
where recall from Section 4 that $X_{\phi}$ denotes the skew symmetric matrix whose entries below the diagonal are the entries of the vector $\phi$.
When $n=2,3$, we have $\det(I+X_{\phi})=1+\lVert\phi\rVert^{2}$ and $\det(I+\gamma X_{\phi})=1+\gamma^{2}\lVert\phi\rVert^{2}$, hence
$\textnormal{KL}(\mathcal{C}_{n}(M,\kappa)\;\|\;U_{SO(n)})=\int_{\mathbb{R}^{d}}\frac{2^{\frac{3n(n-1)}{4}}V_{n}^{-1}\gamma^{d}}{(1+\gamma^{2}\lVert\phi\rVert^{2})^{n-1}}\log\Big{(}\frac{\gamma^{d}(1+\lVert\phi\rVert^{2})^{n-1}}{(1+\gamma^{2}\lVert\phi\rVert^{2})^{n-1}}\Big{)}\,\lambda^{d}(d\phi).$
Recall that by [49],
$V_{n}=2^{\frac{3n(n-1)}{4}}\prod_{i=2}^{n}\frac{\pi^{(i-1)/2}\Gamma((i-1)/2)}{\Gamma(i-1)}.$
Hence $\textnormal{KL}(\mathcal{C}_{n}(M,\kappa)\;\|\;U_{SO(n)})$ is precisely the KL divergence of $\operatorname{Cauchy}_{d}(0,\frac{1}{\gamma^{2}}I)$ from $\operatorname{Cauchy}_{d}(0,I)$ for $n=2$ or $n=3$. By the expressions for the KL divergence between Cauchy distributions in [10], we have
$\textnormal{KL}(\operatorname{Cauchy}_{1}(0,\tfrac{1}{\gamma^{2}}I)\;\|\;\operatorname{Cauchy}_{1}(0,I))=\log\bigg{(}\frac{(\gamma+1)^{2}}{4\gamma}\bigg{)}$
and
$\textnormal{KL}(\operatorname{Cauchy}_{3}(0,\tfrac{1}{\gamma^{2}}I_{3})\;\|\;\operatorname{Cauchy}_{3}(0,I_{3}))=\log\bigg{(}\frac{(\gamma+1)^{4}}{16\gamma}\bigg{)}+\frac{2(1-\gamma)}{1+\gamma}.$
Since $\gamma=\frac{1+\kappa}{1-\kappa}$ we have
$\textnormal{KL}(\mathcal{C}_{2}(M,\kappa)\;\|\;U_{SO(2)})=-\log(1-\kappa^{2})$
and
$\textnormal{KL}(\mathcal{C}_{3}(M,\kappa)\;\|\;U_{SO(3)})=-\log(1-\kappa^{2})-2\log(1-\kappa)-2\kappa.$
Appendix D Experimental Details
All experiments were run on a NVIDIA Tesla K80 GPU with two Intel Xeon E5-2637 CPUs with 64GB of memory. Hyperparameter tuning was done using Weights & Biases [4] to perform a grid search over the hyperparameter configurations described in Table 3. The neural sheaf diffusion (NSD) models from [6] were implemented using the code available at https://github.com/twitter-research/neural-sheaf-diffusion. Our code is available at https://github.com/patrick-gillespie/bsnn.
Table 3: Hyperparameter configurations for WebKB experiments.
Hyperparameter Values
Hidden channels {8, 32}
Stalk dim $d$2, 3, 4, 5}
Layers {2, 3, 4, 5}
Dropout {0.0, 0.3, 0.6}
Learning rate 0.01
Activation ELU
Weight decay 5e-4
Patience 200
Max training epochs 500
Optimiser Adam
Num. models in ensemble 3 (BSNN), N/A (NSD)
D.1 WebKB Dataset Descriptions
The Texas, Cornell, and Wisconsin datasets are web page datasets for the computer science departments of the respective universities. Each dataset is a graph in which nodes represent web pages and edges represent hyperlinks between them. In all datasets the node features are $1703$--dimensional bag-of-word representations of the web pages, and each node belongs to one of the five classes: student, project, course, staff, and faculty. The Texas dataset contains 183 nodes and 325 edges. The Cornell dataset contains 183 nodes and 298 edges. The Wisconsin dataset contains 251 nodes and 515 edges.